# Existence and Uniqueness Results for a Mean-field Game of Optimal Investment

### v2

@@
Abstract.
@@
a University of Parma, Department of Economics and Management, Via J. F. Kennedy 6, 43125 Parma (Italy).
@@
1 E-mail: alessandro.calvia@unipr.it.
@@ 1. Introduction
In this paper, we provide existence, uniqueness, and characterization results for the equilibrium of a mean-field model of optimal investment with competition à la Cournot. In the absence of interventions, the production capacity of the representative company evolves stochastically over time as a geometric Brownian motion, and its level can be increased through investment, subject to quadratic costs. The representative company discounts profits and costs at a constant rate and aims to maximize total expected profits from production, net of investment costs. Instantaneous profits depend linearly on the company’s production capacity (thus, production occurs at full capacity) and on the time-dependent price of the produced good. The mean-field equilibrium investment and average production processes $(\widehat{\bf{u}},\widehat{q})$ are such that expected total net profits are maximized and, assuming an isoelastic inverse demand function, the price is given in terms of a nonlinear function of the average optimally controlled production at each time (see also Achdou et al. [2] and Section 2.1 below). We are able to prove the existence and uniqueness of the equilibrium pair $(\widehat{\bf{u}},\widehat{q})$ when the problem’s time horizon is finite or infinite. These equilibria are characterized as the unique solution to a nonstandard system of forward-backward ordinary differential equations. Furthermore, we demonstrate that the existence and uniqueness carry over to the deterministic counterpart of our model, essentially using the same techniques as in the stochastic case.
@@
It is important to note that although the company’s optimal control problem for a fixed average production trajectory $q$ is linear-quadratic, the overall mean-field problem is not of linear-quadratic type, due to the nonlinear dependence of the net profit functional on $q$. In this regard, our results differ from those presented in Bensoussan et al. [3], Delarue and Tchuendom [12], Tchuendom [20], among others, which focus on linear-quadratic mean-field games.
While the three-step approach previously outlined works for the case where $T<+\infty$ as well as the case where $T=+\infty$, different technical arguments are used to prove the existence and uniqueness of the equilibrium in the finite and infinite horizon cases. As a matter of fact, the integral equation that uniquely characterizes the equilibrium average production $\widehat{q}$ does not fall into the standard theory of integro-differential equations, as we have an initial-value problem with the integral being backward in time, and this requires a careful ad hoc analysis. Specifically, the proof of the existence of a solution relies on a priori estimates and Schauder’s fixed point theorem when $T<+\infty$, whereas it utilizes a priori estimates and the Frechet-Kolmorogorov ${\mathrm{L}}^{p}$ compactness theorem in the case of $T=+\infty$. Uniqueness is then established in both cases through suitable contradiction arguments that exploit the properties of mean-field equilibria. Notably, when $T=+\infty$, we can also prove that the time-dependent equilibrium $\widehat{q}$ monotonically converges to the stationary average production level. The latter is obtained as the unique constant solution to the integral equation that uniquely determines $\widehat{q}$.
@@
Games of optimal investment in both stochastic and deterministic settings are extensively covered in the Economics literature, and a comprehensive overview of models and results can be found in Vives [21]. Specifically, mean-field problems with Cournot competition have garnered interest in recent literature; see Chan and Sircar [10] for an insightful overview. Graber and Bensoussan [15] study the existence (and, under certain conditions, uniqueness) of a solution to a system of partial differential equations (PDEs) (namely, the Hamilton-Jacobi-Bellman (HJB) and Fokker-Planck equations) associated with a mean-field game involving Bertrand and Cournot competition among a continuum of players. In Graber and Sircar [14], existence and uniqueness of the master equation associated with a mean-field game of controls with absorption are proven, while Chan and Sircar [9] delve into dynamic mean-field games with exhaustible capacities and interactions akin to Cournot and Bertrand competitions. An optimal transport perspective on Cournot-Nash equilibria is explored in Acciaio et al. [1]. Finally, Cao et al. [6] considers stationary discounted and ergodic mean-field games of singular controls motivated by irreversible investment and provide existence and uniqueness results, as well as relations across the two classes of considered problems.
To the best of our knowledge, the existence and uniqueness of the (nonstationary) mean-field equilibrium for a mean-field model of optimal investment with an isoelastic demand function, as discussed in this paper, is presented here for the first time. It is also worth noting that some of our results – in particular the existence and uniqueness ones – carry over also to the case of a more general instantaneous profit function and more general mean-field interaction, upon requiring suitable monotonicity conditions on the problem’s data. We refer to Section 2.1 below for further details.
@@ 1.1. Notation
  • Throughout the paper the set $\mathbb{N}$ denotes the set of natural integers without the zero element, i.e., =$\mathbb{N}=\{1,2,\dots\}$, while $\mathbb{R}$ denotes the set of real numbers. Whenever $T=+\infty$, the notation $[0,T]$ indicates the interval $[0,+\infty).$
  • For any $p\geq 1$, any measure space $(E,\mathscr{E},\mu)$, and any interval $I\subseteq\mathbb{R}$, we indicate by ${\mathrm{L}}^{p}((E,\mathscr{E},\mu);I)$ the set of all functions with values in $I$ that are $p$-integrable with respect to $\mu$. If $E\subseteq\mathbb{R}$, we take $\mathscr{L}(E)$ as the Lebesgue $\sigma$-algebra and as $\mu$ the Lebesgue measure, which is denoted by $\mathrm{Leb}$, and we simply write ${\mathrm{L}}^{p}(E;I)$. When computing integrals with respect to this measure, we simply write ${\mathrm{d}}x$ instead of $\mathrm{Leb}({\mathrm{d}}x)$.
@@
  • Given two intervals $I,J\subseteq\mathbb{R}$, we denote by ${\mathrm{C}}(I;J)$ the set of all continuous functions from $I$ to $J$, endowed with the usual sup-norm. The notation ${\mathrm{C}}^{p}(I;J)$, $p\in\mathbb{N}\cup\{+\infty\}$, denotes the set of functions from $I$ to $J$ that are continuously differentiable $p$ times.
@@ 2. The stochastic model
Let $(\Omega,{\mathcal{F}},{\mathbb{F}}\coloneqq({\mathcal{F}}_{s})_{s\geq 0},{\mathbb{P}})$ be a complete filtered probability space, with ${\mathbb{F}}$ satisfying the usual assumptions, supporting an ${\mathbb{F}}$-adapted standard Brownian motion $B$ and an independent ${\mathcal{F}}_{0}$-measurable random variable $\xi$. Throughout the paper, $T\in(0,+\infty)\cup\{+\infty\}$ denotes a finite or infinite time horizon. In what follows, whenever $T=+\infty$ the notation $[0,T]$ indicates the interval $[0,+\infty)$.
@@
where $\delta,\sigma>0$ are given coefficients, and the control process ${\mathbf{u}}\coloneqq(u_{s})_{s\in[0,T]}$ is chosen in either of the following two classes of admissible controls: if $T<+\infty$,
[…]
if, instead, $T=+\infty$,
[…]
Whenever necessary, to stress the dependence of the solution to (2.1) on the initial condition $\xi$ and on the control ${\mathbf{u}}$, we denote it by $X^{\xi,{\mathbf{u}}}$.
We observe that the solution to SDE (2.1) can be written explicitly (cf. [17, Problem 5.6.15]), for each ${\mathbf{u}}\in{\mathcal{U}}_{T}$, as
@@
In the mean-field game studied in this paper, the process $X$ describes the private state of a representative player. In particular, it models the evolution of her/his production capacity, which depreciates at a rate $\delta$ and can be increased by choosing the investment rate ${\mathbf{u}}\in{\mathcal{U}}_{T}$. Note that the statistical distribution of the state of the other players in the economy does not influence the production capacity level of the representative player.
@@
The assumption above ensures that, for any ${\mathbf{u}}\in{\mathcal{U}}_{T}$, $X_{s}^{\xi,{\mathbf{u}}}\geq 0$, ${\mathbb{P}}$-a.s., for all $s\in[0,T]$. Thus, the production capacity level of the representative agent is never negative. Moreover, it grants us the following result, which will be used later on.
@@
Under Assumption 2.1 and for any ${\mathbf{u}}\in{\mathcal{U}}_{T}$, the unique solution to SDE (2.1) has finite first moment, given by
@@
Here, $\rho>0$ is a discount factor, $\beta>0$ is a fixed parameter, and $q=(q_{s})_{s\in[0,T]}$ is a given deterministic measurable function. At equilibrium, the function $q$ will identify with the average production capacity of the whole population of agents. This is formalized in the next definition.
@@
A pair $(\widehat{\mathbf{u}},\widehat{q})$, where $\widehat{\mathbf{u}}\in{\mathcal{U}}_{T}$ and $\widehat{q}\colon[0,T]\to(0,+\infty)$ is a measurable function, is an equilibrium of the mean-field game if
@@ 2.1. The economic foundation
$(x_{1},\dots,x_{n})\mapsto x_{i}\biggl{(}\frac{1}{N}\sum_{j=1}^{N}x^{j}\biggr{)}^{-\beta},$ (2.10)
@@
where $\gamma$ is the elasticity of substitution, $\pi$ is the price set by firm $i$ and $P=\left(\frac{1}{N}\sum_{j=1}^{N}\pi_{j}^{1-\gamma}\right)^{1/(1-\gamma)}$ is a price index. This choice of the demand function is standard in economics and is justified by the Spence-Dixit-Stiglitz preferences (see, e.g., Dixit and Stiglitz [13]).
@@
$\pi_{i}\mapsto\left\{\pi_{i}\left(\frac{\pi_{i}}{P}\right)^{-\gamma}-\frac{1}{x_{i}^{\delta}}\left(\frac{\pi_{i}}{P}\right)^{-\gamma}\right\},$
where $x_{i}>0$ is the productivity (or size) of firm $i$ and, hence, costs decrease with respect to productivity, according to the factor $x_{i}^{-\delta}$, where $\delta>0$ is a given parameter. We can consider each firm production capacity as a proxy for $x_{i}$ and, thus, this setting is coherent with the model introduced above.
@@
$\frac{1}{\gamma-1}\left(\frac{\gamma}{\gamma-1}\right)^{-\gamma}x_{i}^{\delta(\gamma-1)}P^{\gamma}\,,$ (2.11)
which is attained setting the price $\pi_{i}=\frac{\gamma}{(\gamma-1)}x_{i}^{-\delta}$.
This entails that $P=\frac{\gamma}{(\gamma-1)}\left(\frac{1}{N}\sum_{j=1}^{N}x_{j}^{\delta(\gamma-1)}\right)^{-1/(\gamma-1)}$, and hence, substituting this expression into (2.11), we get that the profit is
$\frac{1}{\gamma-1}x_{i}^{\delta(\gamma-1)}\left(\frac{1}{N}\sum_{j=1}^{N}x_{j}^{\delta(\gamma-1)}\right)^{-\frac{\gamma}{\gamma-1}}\,.$ (2.12)
Therefore, setting $\beta=\frac{\gamma}{\gamma-1}$, $\delta=\frac{1}{\gamma-1}$, and $p=\frac{1}{N}\sum_{j=1}^{N}x_{j}$, we get the profit function (2.10) appearing in our model, except for the constant $\frac{1}{\gamma-1}$ which does not affect the optimization problem.

### v3

@@
Existence and Uniqueness Results
for a Mean-field Game of Optimal Investment
Alessandro Calviaa,1 , Salvatore Federicob,2 , Giorgio Ferraric,3 and Fausto Gozzid,4
Sommario.
@@
a Department of Mathematics, Politecnico di Milano Piazza Leonardo da Vinci, 32, 20133 Milano (Italy).
@@
1 E-mail: alessandro.calvia@polimi.it.
@@ 1. Introduction
[…]
@@
It is important to note that although the company’s optimal control problem for a fixed average production trajectory $q$ is linear-quadratic, the overall mean-field problem is not of linear-quadratic type, due to the nonlinear dependence of the net profit functional on $q$. In this regard, our results differ from those presented in Bensoussan et al. [5], Delarue and Tchuendom [15], Tchuendom [29], among others, which focus on linear-quadratic mean-field games. It is also worth stressing that some of our results – particularly the existence and uniqueness results – extend to the case of an instantaneous profit function that is more general than the isoelastic demand function considered in this paper, as well as to more general mean-field interactions, provided that suitable monotonicity conditions are imposed on the problem’s data. We refer to Section 2.1 for further details.
For our existence and uniqueness results, we do not rely on the study of the PDE system associated with the mean-field game (the Hamilton-Jacobi-Bellman equation and the Fokker-Planck equation), nor on the related master equation; we refer to the subsequent literature review for a discussion of related studies offering a PDE perspective on the problem. Instead, we provide an ad hoc analysis that hinges on the linear-quadratic optimization problem at hand and on the nonlinear mean-field interaction term under consideration.
This leads to a characterization of the unique equilibrium, both in the cases $T<+\infty$ and $T=+\infty$, through a nonlinear integral equation for the average production $\widehat{q}$. However, different technical arguments are used to prove existence and uniqueness in the finite- and infinite-horizon settings. Indeed, the integral equation that uniquely characterizes the equilibrium average production $\widehat{q}$ does not fall within the standard theory of integro-differential equations (see, e.g. Gripenberg et al. [23], Lakshmikantham and Rama Mohana Rao [26]), since we face an initial-value problem with a backward-in-time integral term, which requires a careful specific analysis.
More in detail, the proof of existence relies on a priori estimates and Schauder’s fixed-point theorem when $T<+\infty$, whereas in the case $T=+\infty$ it relies on a priori estimates and the Fréchet–Kolmogorov compactness theorem in $L^{p}$. Uniqueness is established in both cases through suitable contradiction arguments that exploit structural properties of mean-field equilibria. Notably, when $T=+\infty$, we also prove that the time-dependent equilibrium $\widehat{q}$ converges monotonically to the stationary average production level, which is obtained as the unique constant solution of the integral equation characterizing $\widehat{q}$.
@@
Games of optimal investment into production capacity, both in stochastic and deterministic settings, are extensively studied in the economics literature, and a comprehensive overview of models and results can be found in Vives [30]. Specifically, mean-field problems with Cournot competition have attracted considerable attention in the recent literature; see Chan and Sircar [13] for an insightful overview. We now provide a non-exhaustive review of papers dealing with mean-field models of investment in capacity that are relevant for our setting.
Graber and Bensoussan [20] study the existence (and, under certain conditions, uniqueness) of solutions to a system of partial differential equations (PDEs)—namely, the Hamilton–Jacobi–Bellman (HJB) and Fokker–Planck equations—associated with a mean-field game involving Bertrand and Cournot competition among a continuum of players. In Graber and Sircar [21], existence and uniqueness of the master equation associated with a mean-field game of controls with absorption are established, while Chan and Sircar [12] analyze dynamic mean-field games with exhaustible capacities and interactions akin to Cournot and Bertrand competition. An optimal transport perspective on Cournot–Nash equilibria is explored in Acciaio et al. [1].
The existence of classical solutions to an extended mean-field game arising in the study of exhaustible resource production under market competition is addressed in Graber et al. [22], where uniqueness of the equilibrium is shown to hold if the Hamiltonian satisfies certain smallness and uniform convexity conditions. Economically, this corresponds to imposing an upper bound on the degree of substitutability among competing firms.
Camilli et al. [8] also consider Cournot mean-field games of controls for the production of an exhaustible resource by a continuum of producers. They establish uniqueness of the solution under general assumptions on the price function and prove convergence of a learning algorithm, which in turn yields existence of a solution to the mean-field game system.
In Escribe et al. [17], a stylized model for investment in renewable power plants under long-term uncertainty—arising from weather conditions, demand fluctuations, future fuel prices, and average national weather patterns—is considered. It is shown that the $N$-player formulation of the problem admits a Nash equilibrium and that any sequence of Nash equilibria of the $N$-player game converges to the unique solution of the corresponding mean-field game.
Investment in renewable energy production is also addressed in Alasseur et al. [4], which proposes a model with a large number of small competitive producers to study the effect of subsidies on the aggregate level of capacity, taking into account a cannibalization effect. The analysis is based on the master equation, and explicit formulas for the long-run equilibria are obtained.
Related price formation models, in the context of large systems of small agents trading a commodity or a security, are studied in Fujii and Takahashi [18], Gomes and Saúde [19] using mean-field techniques.
Regarding irreversible investment, we refer to Cao et al. [9] and Aïd et al. [3]. Cao et al. [9] consider stationary discounted and ergodic mean-field games of singular controls motivated by irreversible investment and provide existence and uniqueness results, as well as connections between the two classes of problems. Applications to commodity markets and the role of regime switches in equilibrium formation are investigated in Aïd et al. [3], which studies a mean-field model of firms competing à la Cournot and irreversibly investing in production.
@@ 1.1. Notation
  • Throughout the paper the set $\mathbb{N}$ denotes the set of natural integers without the zero element, i.e., $\mathbb{N}=\{1,2,\dots\}$, while $\mathbb{R}$ denotes the set of real numbers. Whenever $T=+\infty$, the notation $[0,T]$ indicates the interval $0,+\infty).$
  • For any $p\geq 1$, any measure space $(E,\mathscr{E},\mu)$, and any interval $I\subseteq\mathbb{R}$, we indicate by ${\mathrm{L}}^{p}((E,\mathscr{E},\mu);I)$ the set of all functions with values in $I$ that are $p$-integrable with respect to $\mu$. If $E\subseteq\mathbb{R}$ and it is Lebesgue-measurable, we take $\mathscr{E}=\mathscr{L}(E)$ as the Lebesgue $\sigma$-algebra and as $\mu$ the Lebesgue measure, which is denoted by $\mathrm{Leb}$, and we simply write ${\mathrm{L}}^{p}(E;I)$. When computing integrals with respect to this measure, we simply write ${\mathrm{d}}x$ instead of $\mathrm{Leb}({\mathrm{d}}x)$.
@@
  • Given two intervals $I,J\subseteq\mathbb{R}$, we denote by ${\mathrm{C}}(I;J)$ the set of all continuous functions from $I$ to $J$. If $I$ is compact, we endow this set with the usual sup-norm, which makes it a Banach space. The notation ${\mathrm{C}}^{p}(I;J)$, $p\in\mathbb{N}\cup\{+\infty\}$, denotes the set of functions from $I$ to $J$ that are continuously differentiable $p$ times.
@@ 2. The stochastic model
Let $(\Omega,{\mathcal{F}},{\mathbb{F}}\coloneqq({\mathcal{F}}_{s})_{s\geq 0},{\mathbb{P}})$ be a complete filtered probability space, with $\mathbb{F}$} satisfying the usual assumptions, supporting an $\mathbb{F}$}-adapted standard Brownian motion $B$ and an independent ${\mathcal{F}}_{0}$-measurable random variable $\xi$. Throughout the paper, $T\in(0,+\infty)\cup\{+\infty\}$ denotes a finite or infinite time horizon.
@@
where $\delta,\sigma>0$ are given coefficients, and ${\mathbf{u}}\in{\mathrm{L}}^{1}(\Omega\times[0,T];[0,+\infty))$, i.e., $\mathbf{u}$} is a non-negative, $({\mathcal{F}}_{s})_{s\in[0,T]}$-progressively measurable process satisfying ${\mathbb{E}}\left[\int_{0}^{t}u_{s}\,{\mathrm{d}}s\right]<+\infty$, for all $t\in[0,T]$. Process $\mathbf{u}$} is the control, which we will choose in a suitable class of admissible processes, that will be specified below. Whenever necessary, to stress the dependence of the solution to (2.1) on the initial condition $\xi$ and on the control $\mathbf{u}$}, we denote it by $X^{\xi,{\mathbf{u}}}$.
We observe that SDE (2.1) admits a unique strong solution, which can be written explicitly (cf. [25, Problem 5.6.15]), for each ${\mathbf{u}}\in{\mathrm{L}}^{1}(\Omega\times[0,T];[0,+\infty))$, as
@@
In the mean-field game studied in this paper, the process $X$ describes the private state of a representative player. In particular, it models the evolution of her/his production capacity, which depreciates at a rate $\delta$ and can be increased by choosing the investment rate $\mathbf{u}$}. Since we require process $\mathbf{u}$} to be non-negative, we are in a setting where investment actions are irreversible. From a mathematical point of view, this ensures useful properties of process $X$, which will be discussed below and used throughout the paper. Note also that the statistical distribution of the state of the other players in the economy does not influence the production capacity level of the representative player.
@@
The assumption above ensures that, for any ${\mathbf{u}}\in{\mathrm{L}}^{1}(\Omega\times[0,T];[0,+\infty))$, $X_{s}^{\xi,{\mathbf{u}}}>0$, $\mathbb{P}$}-a.s., for all $s\in[0,T]$. Thus, the production capacity level of the representative agent is never negative. Moreover, it grants us the following result, which will be used later on.
@@
Under Assumption 2.1 and for any ${\mathbf{u}}\in{\mathrm{L}}^{1}(\Omega\times[0,T];[0,+\infty))$, the unique solution to SDE (2.1) has finite first moment, given by
@@
Here, $\rho>0$ is a discount factor, $\beta>0$ is a fixed parameter, and $q=(q_{s})_{s\in[0,T]}$ is a given deterministic measurable function. The control process $\mathbf{u}$} is chosen in either of the following two classes of admissible controls: if $T<+\infty$,
${\mathcal{U}}_{T}\coloneqq$ $\biggl\{{\mathbf{u}}=(u_{s})_{s\in[0,T]}\text{ s.t. }{\mathbf{u}}\colon\Omega\times[0,T]\to[0,+\infty)\text{ is }({\mathcal{F}}_{s})_{s\in[0,T]}\text{-progressively measurable and }$
if, instead, $T=+\infty$,
${\mathcal{U}}_{\infty}\coloneqq$ $\biggl\{{\mathbf{u}}=(u_{s})_{s\geq 0}\text{ s.t. }{\mathbf{u}}\colon\Omega\times[0,\infty)\to[0,+\infty)\text{ is }{\mathbb{F}}\text{-progressively measurable,}$
At equilibrium, the function $q$ will identify with the average production capacity of the whole population of agents. This is formalized in the next definition.
@@
[…]
@@
From SDE (2.1), the expression of the expected net profit functional in (2.5), and the definition of equilibrium given above, we see that the mean-field interaction between firms appears only in the functional to be maximized, and not also in the state equation. We also emphasize that the interaction is only through the average production capacity, thus our problem is set in a Cournot-like competition framework and, as we are going to see, the optimal control for the representative firm is going to depend on this quantitity.
The very specific nature of our problem allows us to tackle it with ad hoc arguments, as we outline below, rather than using either the standard analytic approach (see, e.g., [10]), where a system of coupled Hamilton-Jacobi-Bellman and Kolmogorov-Fokker-Plank equations is studied, or the standard probabilistic approach (see, e.g., [11]), where a system of Forward Backward Stochastic Differential Equations is considered.
@@ 2.1. The economic foundation
$(x_{1},\dots,x_{n})\mapsto x_{i}\biggl(\frac{1}{N}\sum_{j=1}^{N}{x_{j}}\biggr)^{-\beta},$ (2.10)
@@
where $\gamma$ is the elasticity of substitution, $\pi_{i}$ is the price set by firm $i$ and $P=\left(\frac{1}{N}\sum_{j=1}^{N}\pi_{j}^{1-\gamma}\right)^{1/(1-\gamma)}$ is a price index. This choice of the demand function is standard in economics and is justified by the Spence-Dixit-Stiglitz preferences (see, e.g., Dixit and Stiglitz [16]).
@@
$\pi_{i}\mapsto\left\{\pi_{i}\left(\frac{\pi_{i}}{P}\right)^{-\gamma}-\frac{1}{x_{i}^{{\eta}}}\left(\frac{\pi_{i}}{P}\right)^{-\gamma}\right\},$
where $x_{i}>0$ is the productivity (or size) of firm $i$ and, hence, costs decrease with respect to productivity, according to the factor $x_{i}^{-{\eta}}$, where $\eta>0$ is a given parameter. We can consider each firm production capacity as a proxy for $x_{i}$ and, thus, this setting is coherent with the model introduced above.
@@
$\frac{1}{\gamma-1}\left(\frac{\gamma}{\gamma-1}\right)^{-\gamma}x_{i}^{{\eta}(\gamma-1)}P^{\gamma}\,$, (2.11)
which is attained setting the price $\pi_{i}=\frac{\gamma}{(\gamma-1)}x_{i}^{-{\eta}}$.
This entails that $P=\frac{\gamma}{(\gamma-1)}\left(\frac{1}{N}\sum_{j=1}^{N}x_{j}^{{\eta}(\gamma-1)}\right)^{-1/(\gamma-1)}$, and hence, substituting this expression into (2.11), we get that the profit is
$\frac{1}{\gamma-1}x_{i}^{{\eta}(\gamma-1)}\left(\frac{1}{N}\sum_{j=1}^{N}x_{j}^{{\eta}(\gamma-1)}\right)^{-\frac{\gamma}{\gamma-1}}\,.$ (2.12)
Therefore, setting $\beta=\frac{\gamma}{\gamma-1}$, $\eta=\frac{1}{\gamma-1}$, we get the profit function (2.10), except for the constant $\frac{1}{\gamma-1}$ which does not affect the optimization problem. We emphasize once more that (2.10) justifies the form of the profit function $(x,p)\mapsto xp^{-\beta}$ appearing in (2.5). Indeed, we formally substitute the production capacity $x_{i}$ of firm $i$ with the production capacity $x$ of the representative firm, and we replace $\frac{1}{N}\sum_{j=1}^{N}x_{j}$ with variable $p$ which, at equilibrium, will coincide with the average production capacity of the mean-field game.